Left Direct Gregorian Technical Observing with IRTC: UT 2009 October 10

Observer: JHill(LBTO), DMiller (Tucson)
Telescope Operator: Aaron Ceranski
Instrument Support:TShih
Telescope Support: JRousseau
SW support: (Tucson)

Summary

Poor (1-1.5 arcsec) seeing at the beginning, but more stable than last night. We remeasured the position of the rotator center on Guider and IRTC. We adjusted LEFTZEROPOINT for IRTC on LDG. We took some guider throughput data on M103.

Closed for clouds from 6:25 until 9:40 UT.

Then spent the latter part of the night taking a pretty good set of AGW transform data on the Stone C astrometric field. We were working through scattered cirrus clouds, but the seeing was 0.5 arcsec on the guider. Telescope worked quite well until the last half hour when things got cranky.

Tested a prototype IRAF script to automatically select stars and collect pointing data. (But, didn't take any significant pointing data this night.)

Preparations

Software folks have given us TCS Build 34 which straightens out the problems reading boolean parameters in Build 33. Build 34 also has the first PSF telemetry.

Marco has given us the finder chart for the Stone C field. We are using it from /home/LBTO/iraf/LBTtools/IRTC/Catalogs.

Details

All Times in UT

Open

01:16 Open Vent Doors

01:18 Open Shutter Doors. T=5.3 degC, D=-16 degC, Wind light. Sky mostly clear with wispy cirrus approaching from SW. But mirrors are still down at 1.2 degC!

01:22 Pointto M5_1539 TRACK. Found star at IE=8 CA=39.

01:32 Pointto M5_1539 gs=5 ACTIVE. Initial seeing is around 1 arcsec.

Pointing Stability Testing

01:52 pointto StoneOagw gs=0 ACTIVE PA=0 - Primary hits XY limit at Y=-2.99.

01:56 Adjust M1 and M2 by +1.0 mm in Y.

01:57 pointto StoneOagw gs=0 ACTIVE PA=0 - found star completely at edge of acquisition field.

01:59 pointto StoneOagw gs=0 ACQUIRE PAM- Adjust *IE=7 CA=54 *

02:02 Drive telescope to zenith (trying to reproduce the LUCIFER pointing issues seen last week)

02:04 pointto StoneOagw gs=0 ACQUIRE PAM- - Preset failed because mirror is down. Apparently too much C00. Lost 10 minutes.

02:12 pointto StoneOagw gs=0 ACQUIRE PAM- Adjust IE=7 CA=51

02:14 Drive to zenith again

02:15 pointto StoneOagw gs=0 ACQUIRE PAM- Adjust IE=7 CA=50 - No sign that there is a pointing stability problem. Only the systematic drift from the temperature of the telescope changing rapidly.

02:18 pointto StoneOagw gs=0 ACTIVE PA=0 - Seeing has improved to 0.8 arcsec at EL=51 deg.

02:26 Both John's and Aaron's GCSGUI are spending lots of the time with bogus image displayed. This is now bad enough that it is a problem.

Remeasure Rotator Center on Guider

19:36 Purposefully Offset CA=40

02:37 readGuideCam -e 75000 in /Repository/AGW_Data/091010guider/ will be guider000002 (a nice round ring)

  • Center of Rotation of AGw#4 Guider: (293,264.5) compared to Jesper's (294,264)
    guider000002.jpg

02:39 pointto StoneOagw TRACK

02:41 takeseq command fails.

IRTC> !irc GetFilter
IRS is not running.
Exit 1

02:50 Aaron restarts irs as root. Lost 10 minutes.

02:52 takeseq 1.0 80 irtc.20091010.025246.fits in /Repositoy/AGW_Data/091010/ irc GetSeqImages 80 failed for no obvious reason. cube 0047 (aka flat) seems to exist. I took a short manual sequence and it worked.

02:57 pointto StoneOagw gs=0 TelM="ACTIVE" PA=0

03:01 Primary mirror panics for Z upper limit on actuator 301. But the telescope didn't move, so it isn't clear how an illegal force got applied. Aaron clears C00 and reraises mirror. Lost 10 minutes.

03:04 Guider is detecting noise at the edge of the field. Perhaps we have the sextractor thesholds too low.

03:06 pointto StoneOagw gs=0 TelM="ACTIVE" - Image appears double after C00 is cleared. Could there be something pushing on the mirror? Maybe the high forces aren't coming from Z22 as we thought. Seeing is still around 0.8 arcsec. Star is nearly centered with IE=7 CA=40

LEFTZEROPOINT Adjustment

03:13 Based on Doug's analysis of last night's data. Change LEFTZEROPOINT from +5.5 to +4.45 degrees Stop and restart PCS. Reset IE=7 CA=40

03:15 pointto StoneOagw gs=0 TelM="ACTIVE" PA=0

03:19 pointto StoneOagw gs=0 TelM="TRACK" PA=0

03:20 pointto M5_0855 TRACK

takeseq 0.02 100

irtc.20091010.032246.fits

03:23 offset 0 10 coord=RACDEC ABS

takeseq 0.02 100

offset 0 0

takeseq 0.02 100

Data in /Repository/IRTC_Data/091010/irtc.20091010.032454.fits

03:28 pointto M5_0855 gs=2 TelM="ACTIVE" PA=INDEF

    • Using IDL to stack the cubes and the FIND routine Doug gets
Image X Y
irtc.20091010_00049_cube.fit 201.643 165.521
irtc.20091010_00050_cube.fit 205.830 69.8244
irtc.20091010_00051_cube.fit 205.3346 167.642
    • angle = atan(49&50) = -2.5106
    • angle = atan(50&51) = 0.4065
    • The -2.5 is definitely a result of tracking error, or because the IRTC images were saturatured

Remeasuring Rotator Center on IRTC

03:40 takeseq 0.01 750 frm=10 while Aaron rotates 0-360

Data in /Repository/IRTC_Data/091010/irtc.20091010.033958.fits cube 0053

  • Measured center at 163.37, 132.37 DLM
  • IRTC image with rotation of 0-360:
    irtc.20091010_00053_stacked.jpg

   irc GetSeqImages 750
   > Problem with command
...Done! (failed: see LOGs/takeseq_033953.log)

03:28 pointto M5_0855 gs=2 TelM="ACTIVE" PA=INDEF

03:45 pointto M5_0855 gs=2 TelM="ACTIVE" PA=INDEF

Another check of LEFTZEROPOINT with dice5

03:49 pointto M5_0855 gs=0 TRACK PA=0

Adjust CA=35

03:57 dice5 off1=7 off2=7 coord=RADEC exp=0.01 Nexp=100 loopN=2

Data in /Repository/IRTC_Data/091010 cubes 54-63. Check date for irtc.*.fits for appropriate header information

AGW Throughput Measurements

04:07 pointto M103_OC gs=0 TelM="ACTIVE" PA=0

Sky is pretty clear, but with an occasional jet contrail drifitng by. We'll call it nearly photometric. Seeing is varying between 0.8 and 1 arcsec.

Data in /Repository/AGW_Data/091010guider

04:12 Stop guiding - not sure about exactly where in M103 we are.

04:13 10 sec guider00003, 20 sec guider 00004 Filter=3 (GCS numbering) r'

04:18 20 sec guider00005-7 Filter=2 (GCS) V

04:20 20 sec guider00008-10 Filter=1 (GCS) or Filter=0 (Agw) B

04:22 offset 10 0 coord=RADEC ABS

04:23 20 sec guider00011-13 Filter=1 (GCS)

04:25 20 sec guider00014-16 Filter=2 (GCS)

04:27 20 sec guider00017-19 Filter=3 (GCS)

04:28 pointto M103_OC gs=0 TelM="ACTIVE" PA=0 (same coords but guider grabbed a different star)

04:32 Stop guiding

04:33 20 sec guider 20-22 Filter=3

04:36 20 sec guider 23-25 Filter=2

04:37 20 sec guider 26-28 Filter=1

04:40 pointto M103_OC gs=0 TelM="ACTIVE" PA=0 (grabbed still a different star, this one much fainter) Guiding seems OK on this faint star.

04:47 Stop guiding. This sequence of offsets is to make a little r-band finder chart to make identifying the previous images easier.

04:48 20 sec guider 29-31 Filter=3

offset 20 0 RADEC ABS

04:50 20 sec guider 32-34 Filter=3

offset -20 0

04:52 20 sec guider 35-37 Filter=3

offset 0 20

04:54 20 sec guider 38-40 Filter=3

offset 20 20

04:56 20 sec guider 41-43 Filter=3

offset -20 20

04:57 20 sec guider 44-46 Filter=3

offset 0 -20

04:59 20 sec guider 47-49 Filter=3

offset 20 -20

05:01 20 sec guider 50-52 Filter=3

offset -20 -20

05:03 20 sec guider 53-55 Filter=3

05:06 PointingLog PTLog20091010directIRTC

Pointing Model Data

05:20 pointto BS9107 TelMode=ACQUIRE gs=0 Found two stars on the edgr of the guider

05:24 pointto M5_0094 TelMode=TRACK gs=0 Could not find a star, even with 120" ptspiral search

05:30 pointto M5_1040 TelMode=TRACK gs=0 Big focus, so had increase exp time. Change m2 focus

05:38 pointto BS9107 TelMode=ACQUIRE gs=0 Set IE=5 CA=58

05:40 pointto BS9107 TelMode=ACTIVE gs=0 Collimate

Seeing 0.6-0.7"

Transform Data for Jesper

05:44 pointto StoneAagw TelMode=ACQUIRE gs=0 Set IE=9 CA=50

05:44 pointto StoneAagw TelMode=ACTIVE gs=0

05:55 pointto StoneAagw TelMode=ACTIVE gs=12

ecl> lpar transtar
      gstarID = 0               pointto-Guide star ID from catalog?
    (sourceID = "StoneAagw")    pointto-Source name from catalog?
          (PA = INDEF)          pointto-Position angle to set to?
      (agwNum = 4)              3 for AGW3, 4 for AGW4
     (exptime = 0.2)            takeseq-Exposure time in seconds?
     (Nimages = 100)            takeseq-Number of images to take in sequence?
     (verbose = yes)            Verbose output to screen?

Data in /Repository/IRTC_Data/091010

06:00 cl < tran_StoneA_00.cl
transtar (12) irtc.20091010.060546.fits cube 64
transtar (16)
transtar (17)
transtar (18)
transtar (14) picked CCD noise over the faint star ???
transtar (12)
transtar (6) irtc.20091010.061732.fits cube 69
transtar (5)
transtar (3)
transtar (12)
transtar (11)

Need to take longer IRTC cubes so there is sufficient time for active optics. Clouds are moving in from the East.

Close

06:25 Close for clouds

Aaron is using this cloudy period to do some remodelling of the to-station. Will we ever be able to observe again?

John has finished the gopoint.cl IRAF script which will take pointing data automatically (assuming that the star lands somewhere in the IRTC field-of-view).

Open

09:40 Reopen as the clouds have parted.

09:42 pointto ACT0280 gs=0 TelM="ACQUIRE" PAM- - found in the corner of the acquisition image. Adjust IE=-1 CA=40

09:47 pointto StoneCagw gs=0 TelM="ACTIVE" PA=0 - no star because John left the colons out of the DEC value in LBTdefault.cat.

09:54 pointto StoneCagw gs=0 TelM="ACTIVE" PA=0 - Now we are at the correct coordinates, but still didn't find a star.

9:57 pointto ACT0210 gs=0 TelM="ACQUIRE" PAM- Found star off the right edge of the acq image. Now it seems that we have had a "pointing jump".

10:00 pointto StoneCagw gs=0 TelM="ACTIVE" PA=0 - IE=-2 CA=58

10:03 Clear the Secondary Offload corrections from M1. Made a boatload of coma as expected.

10:08 offset 0 0 to offload secondary

10:10 pointto StoneCagw gs=0 TelM="ACTIVE" PA=0 - IE=-2 CA=45 The circumstantial evidence is that something is causing PSF to send in Secondary Offload commands that are not pointing-free. Is this happening at the end of presets where no active optics corrections were applied?

10:13 pointto StoneCagw gs=16 TelM="ACTIVE" PA=0 - (John read the finder chart wrong.)

AGW Transformation Data on Stone C field

These cubes are being written to /Repository/IRTC_Data/default/ at the moment.

The single IRTC images with header information to go with the cubes are written in /Repository/.

MOVED data to /Repository/IRTC_Data/091010 DLM

The logbook.1 file with getdat info is in /home/agwuser/agw-control on lbtmu105.

The position angle is indicated in the name of the script.

10:15 pointto StoneCagw gs=18 TelM="ACTIVE" PA=0 - Seeing is 0.5 arcsec on the guider. Clouds are still threatening.

10:21 cl < tran_StoneC_00.cl
transtar (18) irtc.20091010.102215.fits pt_00087_cube.fit
transtar (10) irtc.20091010.102420.fits pt_00088_cube.fit
transtar (14) pt_00089_cube.fit
transtar (15) pt_00090_cube.fit
transtar (12) irtc.20091010.103037.fits pt_00091_cube.fit
transtar (18) pt_00092_cube.fit

10:33 cl < tran_StoneC_90.cl
transtar (21) irtc.20091010.103518.fits pt_00093_cube.fit
transtar (20) pt_00094_cube.fit

10:38 cl < tran_StoneC_180.cl
transtar (13) pt_00095_cube.fit
transtar (16) pt_00096_cube.fit
transtar (17) pt_00097_cube.fit
transtar (7) pt_00098_cube.fit

10:46 cl < tran_StoneC_225.cl
transtar (13) pt_00099_cube.fit
transtar (16) pt_00100_cube.fit
transtar (17) pt_00101_cube.fit
transtar (7) pt_00102_cube.fit
transtar (4) pt_00103_cube.fit
transtar (5) pt_00104_cube.fit
transtar (9) irtc.20091010.110056.fits pt_00105_cube.fit
transtar (13) irtc.20091010.110312.fits pt_00106_cube

Note that these stars are at least a magnitude brighter than we need when the seeing is 0.5 arcsec. (and the clouds are not too heavy)

11:04 cl < tran_StoneC_270.cl
transtar (9) pt_00107_cube.fit
transtar (10) irtc.20091010.110849.fits pt_00108_cube.fit
transtar (3)
transtar (7) irtc.20091010.111307.fits pt_00110_cube.fit
transtar (4)
transtar (5) irtc.20091010.111719.fits
transtar (14) irtc.20091010.111932.fits pt_00113_cube.fit

Clouds are getting heavy again.

11:20 cl < tran_StoneC_315.cl
transtar (9)
transtar (10)
transtar (3) irtc.20091010.112647.fits pt_00116_cube.fit
transtar (14)
transtar (15)
transtar (12)

11:35 cl < tran_StoneC_190.cl
transtar (13) irtc.20091010.113645.fits
transtar (16)
transtar (17)
transtar (7)
transtar (4)
transtar (8) irtc.20091010.114705.fits

11:48 cl < tran_StoneC_200.cl
transtar (13):AGw command status: timed out waiting for probe to reach destination
transtar (16) Couldn't move probe
transtar (17) Couldn't move probe
transtar (7) Couldn't move the probe
transtar (8)
transtar (4)
transtar (5)

11:54 Try to home the motors, as AGW seems to be wedged. Didn't work.

11:56 Stop and restart AGW4. Then it homes OK.

11:59 cl < tran_StoneC_200.cl
transtar (13) Didn't get valid CCD images.
transtar (16)

12:01 Stop and restart GCS.

12:02 cl < tran_StoneC_200.cl
transtar (13) ERROR: cannot open `uparm$pipe1211' for writing

12:03 Stop and restart IRAF

12:05 cl < tran_StoneC_200.cl
transtar (13) Didn't find any star

12:07 pointto ACT0210 gs=0 TelM="ACQUIRE" PAM- Adjust IE=-19 CA=52

12:11 cl < tran_StoneC_200.cl
transtar (13) irtc.20091010.121208.fits pt_00132_cube.fit (image isn't well collimated)
transtar (16)
transtar (17) irtc.20091010.121611.fits pt_00134_cube.fit
transtar (7) irtc.20091010.121818.fits
transtar (8) irtc.20091010.122025.fits pt_00136_cube.fit
transtar (4)
transtar (5) pt_00138_cube.fit

12:30 Stacked all of the Transform cubes for Jesper: Cube numbers 64 through 138. NOTE there are some empty images ( about 70-86 and 126-131...be sure and check)

Tests of the Auto-Pointing script

12:25 turn pointing logging back on

12:26 gopoint 1 ploop+

Gopoint worked pretty well, except I had mistakenly been presetting in ACTIVE mode. This turned out to be useful for sucking in bright stars from outside the acquisition field, so we are adopting it as a feature. (Need to fix the script tomorrow.)

Close

12:48 Close for impending dawn. T=5.2 degC D=-17.5 degC. Wind 6 m/sec. Scattered clouds.

-- DougMiller - 29 Sep 2009

I Attachment Action Size Date Who Comment
Jespers_Email_Transform.pdfpdf Jespers_Email_Transform.pdf manage 28 K 11 Oct 2009 - 02:21 DougMiller Transform Results from Jesper
daq.20091010.gifgif daq.20091010.gif manage 39 K 11 Oct 2009 - 04:54 JohnHill Plot of telescope and mirror temperatures
guider000002.jpgjpg guider000002.jpg manage 24 K 11 Oct 2009 - 03:26 DougMiller Center of Rotation of AGw#3 Guider
irtc.20091010_00053_stacked.jpgjpg irtc.20091010_00053_stacked.jpg manage 41 K 10 Oct 2009 - 04:34 DougMiller IRTC image with rotation of 0-360
Topic revision: r17 - 11 Oct 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback